a 50-foot pole is supported by a 150-foot-long guy wire at its top. find the horizontal distance (on the ground) from the pole to the wire.
Answer by rfer(16322)

150^2-50^2=b^2
22500-2500=20000
sqrt 20000=141.4 ft